OpenStudy (anonymous):

A bank offers two interest account plans. Plan A gives you 6% interest compound annual simple interest. Plan B gives you 13% annual simple interest. You plan to invest $ 2,000 for the next 4 years. Which account earns you the most interest ( in dollars) after 4 years? How much will you have earned?

OpenStudy (ranga):

Plug the numbers into the compound interest formula and the simple interest formula and see which one yields the higher interest.

OpenStudy (ranga):

For Plan A, here is the compound interest formula: \[\Large A = P(1 + \frac{ r }{ n })^{nt}\]A = Amount at maturity P = Principal Amount r = Annual interest rate in decimal n = compounding period (compounded how many times a year) t = years invested Interest earned = A - P For Plan B, here is the simple interest formula: \[\Large I = P \times r \times t \]